OpenStudy (anonymous):

can anyone solve this? 4^(x+7)=6^(x-1)?

OpenStudy (anonymous):

what kind of math is this? Like algebra?

sam (.sam.):

4^(x+7)=6^(x-1) ln(4^(x+7))=ln(6^(x-1)) xln(4)+7ln(4)=ln(6^(x-1)) xln(4)+7ln(4)-xln(6)=-ln(6) x(ln(4)-ln(6))=-7ln(4)-ln(6) x=(-7ln(4)-ln(6))/(ln(4)-ln(6)) x=-28.3522

OpenStudy (anonymous):

oh yes it is algerbra !... and yei i think i finally got it! got the same!! thank u so much .sam!